The directins read: Use a proportion to answer the question.
Question 1.: What percent of 125 is 25?
I did this: 25/125 = x/125 = 125x/125 = 2500/125
x= 20

Answer by checkley77(12844)

25/125=x/100
125x=2500
x=2500/125
x=20%